OpenStudy (anonymous):

magnitude of the complex vector: (pi+i)^54 ---------- (pi-i)^54

OpenStudy (anonymous):

just to be clear its: \[\left| \frac{ (\pi-i)^(54) }{ (\pi-i)^(54) } \right|\]

OpenStudy (nincompoop):

isn't that just 1 ?

OpenStudy (anonymous):

though the exponents are a bit messy there... Anyways, I've tried bringing the power outside the magnitude, then multiplying by the conjugate of the denominator to get it in terms of real numbers, then taking that result and multiplying by the conjugate of the numerator, then taking the magnitude. I ended up with  which was incorrect. and yes the answers 1 but how if I might ask?